### v1

# Exploring Compositional Generalization (in ReCOGS_pos) by Transformers using Restricted Access Sequence Processing (RASP)

Abstract
Humans understand new combinations of words encountered if they are combinations of words recognized from different contexts, an ability called Compositional Generalization. The COGS benchmark (Kim and Linzen, 2020) reports 0% accuracy for Transformer models on some structural generalizations. We use (Weiss et al., 2021)’s Restricted Access Sequence Processing (RASP), a Transformer-equivalent programming language, to prove by construction that a Transformer encoder-decoder can perform the semantically equivalent ReCOGS_pos (Wu et al., 2024) variant of COGS systematically and compositionally: Our RASP model attains 100% semantic exact match111and 100% string exact match on the ReCOGS test set and 100% SEM on all generalization splits except obj_pp_to_subj_pp which gets 92%. Furthermore, our RASP model shows the ReCOGS_pos task does not require a hierarchical or tree-structured solution: we use word-level tokens with an "embedding" layer that tags with possible part of speech222per (Tenney et al., 2019) by layer 0 the part-of-speech could be predicted for most words in Transformers trained on a masked language modeling objective, so we assume an equivalent embedding is learnable., applying just once per encoder pass 19 attention-head compatible flat pattern-matching rules, shown using grammar coverage (Zeller et al., 2023) to be learnable from the training data, plus general prepositional phrase (pp) handling and sentential complement (cp) handling logic, and output the next logical form (LF) token (repeating until the LF is complete). The model does not apply recursive, tree-structured rules like ‘np_det pp np -> np_pp -> np‘, but scores 100% semantic and string exact match on pp recursion, cp recursion using the decoder loop.
1 Introduction
It was long argued that connectionist models (i.e. neural networks) were somehow structurally incapable of compositional generalization (Fodor and Pylyshyn, 1988). [^3] More specific versions of this debate continue, for example re: syntax, one can read (van Schijndel et al., 2019) vs (Goldberg, 2019) or re: hierarchical generalization by Transformers, (Petty and Frank, 2021) vs (Murty et al., 2023). However, large language models based on the Transformer architecture (Vaswani et al., 2017) compose seemingly fluent and novel text and are excellent few or zero shot learners (Brown et al., 2020).
Some observations do contradict that Transformers learn systematic, compositional solutions to problems that generalize [^4] See ”Appendix: Composition and Learning” (9.13), for example structural generalizations in the COGS task and ReCOGS(Wu et al., 2024) variant of the COGS task (Kim and Linzen, 2020) , a benchmark based on extracting semantics (logical form) from the syntax (grammatical form) of synthetic sentences in a simplified subset of English grammar, requiring models trained only on certain grammar examples to generalize to sentences with unseen grammar built up / recombined from parts present in the training examples.
We use (Weiss et al., 2021)’s Restricted Access Sequence Processing (RASP) language that can be compiled to concrete Transformer weights to prove by construction that a Transformer encoder-decoder [^5] We follow (Zhou et al., 2023) who used RASP to analyze auto-regressive decoder-loop cases, not just Transformer encoders as done by RASP author (Weiss et al., 2021). can perform ReCOGS_pos666official variant, closer to COGS than non-positional ReCOGS, as COGS is also positional, and means we can also measure string exact match, not just semantic exact match over the vocabulary and grammar of that task in a systematic, compositional way (length and recursion depth limited) as a rigorous starting point to investigating when Transformers might learn or not actually learn such compositional/systematic solutions. We find a flat, not hierarchical/tree-structured model which lacks any handling for the recursive rules in the grammar (for prepositional phrase recursion and sentential complement recursion) can perform the task at high accuracy, but requires a special rule for avoiding "attraction" errors [^7] These attraction errors are similar to those discussed elsewhere in NLP and psycholinguistics literature on hierarchical vs linear processing by language models and humans, see ”Appendix: Attraction errors” (9.6). where inserted prepositional phrase nouns replace agent/theme/recipient nouns in the logical form by accident. This is our main result and suggests that the ReCOGS task can be performed with high accuracy by Transformers, turning efforts to learnability, and also adds to the literature a caveat on interpreting success on ReCOGS (and COGS888semantically equivalent and see also https://github.com/willy-b/RASP-for-COGS) by noting a hierarchical or tree-structured representation is not necessarily required (contrary to (Kim and Linzen, 2020) and assumption of (Murty et al., 2022)).
Figure 1: Introducing parse trees, logical form, and semantic graphs. Two semantically identical but syntactically distinct (i) sentences (a) "a boy painted the girl" and (b) "the girl was painted by a boy" are shown with (ii) their distinct parse tree (parsed into COGS input grammar), (iii) the string form of their semantics (ReCOGS logical form; differs in indices and ordering), and (c) the graph representation of their logical form (semantic graph [^10] As a convention, in converting ReCOGS logical forms to Semantic Graphs we use the logical form (source, target) index order for directed semantic graph edges (from verb to related entity) EXCEPT for the agent relationship which is from the agent of a verb to the verb (opposite direction from logical form in that case), which gives our semantic graphs of ReCOGS sentences an unambiguous starting point (layout starts from agent) without affecting comparison of the graphs (generated by a consistent rule), see also Figure 3., not different at all between the two examples). Note the (iii) logical forms (LFs) differ by String Exact Match but not (Wu et al., 2024)’s Semantic Exact Match (order and indices do not match but nouns, normalized verbs, and relationships between nouns and verbs are same). Note the "agent", "theme" order in the logical form string is not required to match for Semantic Exact Match. COGS and ReCOGS tasks require extracting the semantics/meaning (c) encoded in LFs (iii) of sentences (i).
Finally, we predict that these "attraction" errors we had to specifically avoid in our RASP model are contributing to the high error rate of the (Wu et al., 2024) baseline Transformer trained from scratch and confirm this is the case.
2 Prior Literature
(Kim and Linzen, 2020) introduce the COmpositional Generalization Challenge based on Semantic Interpretation (COGS) benchmark111111 See Figure 10. and argue that Transformers have low accuracy on the generalization splits (35% overall), especially structural generalization splits where near 0% accuracy is reported, using a 2-layer Encoder-Decoder Transformer (2 layers for Encoder, 2 layers for Decoder).
For another example, (Lake and Baroni, 2023) use a "meta-learning for compositionality" approach with a 3-layer Encoder-Decoder Transformer architecture and achieve what they call "human-like systematic generalization", achieving high scores on everything in the COGS benchmark (>99% on lexical generalizations) EXCEPT the structural generalization splits where they also still score 0% accuracy. However, one notices these networks are shallow compared with those used in successful large-pretrained Transformer models (e.g. 24-layer BERT where compositional parse trees seem to be encoded in its vector space representation (Hewitt and Manning, 2019)), and it is claimed, by e.g. (Csordás et al., 2022) that for compositional operations, like parsing, the depth of the network must be at least the maximum number of compositional operations, e.g. the height of the parse tree for grammar dependent problems. Remarkably, (Petty et al., 2024) finds that increasing the layer depth of the Transformer models (up to 32 layers) does not improve the near 0% accuracy on COGS structural generalization splits like prepositional phrase modification of subject when the network has only seen it on the object during training and also input length/depth generalizations (like pp/cp recursion), perhaps surprising as for the simpler logical inferences problem in (Clark et al., 2020) they observed successful logical inference depth generalization even by Encoder-only Transformers.
Thankfully, (Wu et al., 2024) are able to begin to get traction (low but nonzero accuracy) for the shallow Encoder-Decoder Transformer models on structural generalizations in a modified but semantically equivalent form of the COGS task they call ReCOGS, which we analyze here. They remove redundant symbols, and use Semantic Exact Match instead of Exact Match (see Figure 10).
(Zhou et al., 2023) apply (Weiss et al., 2021)’s RASP language to explain some inconsistent findings regarding generalization and use RASP to predict exactly which cases of generalization come easily to Transformers and which do not. (Zhou et al., 2023) seem to reveal (Weiss et al., 2021) has provided the framework we seek by demonstrating how to apply RASP to Transformer decoders with intermediate steps, and even use it to learn how to modify difficult-to-learn tasks like Parity [^12] See (Strobl et al., 2024) for context from formal language theory, computational complexity, circuit complexity theory, and experimental papers together, providing robust lower and upper bounds on what Transformers can do, including discussion of under what conditions Parity can be solved by Transformers and how whether it can be learned by randomly initialized Transformers under simple training schemes is a different question (general feed-forward neural networks can learn to solve Parity per (Rumelhart et al., 1988)). (Delétang et al., 2023) also. See ”Appendix: Zhou et al 2024 relevance of their long addition experiment to language modeling and note on the Parity task and Transformers” (9.12) and long addition in seemingly incidental ways based on RASP analysis to make them readily learnable by Transformers in a compositional, length generalizing way!
Thus we apply techniques similar to (Zhou et al., 2023) and (Weiss et al., 2021) to ReCOGS to (1) argue Transformers should be capable of performing the task, including the structural generalization splits, with high accuracy, and that the problem is learning not capability and (2) try to understand the prepositional phrase modification related generalization errors (Wu et al., 2024)’s baseline Encoder-Decoder Transformers are making.
3 Data
The ReCOGS (Wu et al., 2024) dataset (ReCOGS with positional indexing variant) based on COGS (Kim and Linzen, 2020) (input sentences/grammar are the same but logical form output are different) was used [^13] https://github.com/frankaging/ReCOGS, with special attention on the structural generalization splits (especially prepositional phrase Object-to-Subject generalization).
The grammar and vocabulary description for COGS English input sentences provided in the utilities associated with the IBM CPG project (Klinger et al., 2024) [^14] https://github.com/IBM/cpg/blob/ c3626b4e03bfc681be2c2a5b23da0b48abe6f570 /src/model/cogs_data.py#L523 were used in designing our RASP solution and analyzing the ways in which this task could be learned (we did not actually use their grammar though, and our RASP solution is flat and non-hierarchical unlike their description of the COGS probabilistic context-free grammar which is hierarchical and recursive). See Figure 10.
4 Model
We used the RASP interpreter of (Weiss et al., 2021). For RASP model design and details see Appendix 9.2. We use word-level tokens for all RASP model results in this paper, [^15] We believe any solution at the word-level can be converted to a character-level token solution (see Appendix 9.3 for proof of concept details on a character level solution not used here). with an "embedding" layer that tags with possible part of speech [^16] Note we follow the (Klinger et al., 2024) description of COGS and include in our RASP vocabulary (part-of-speech or verb-type embedding/mapping) all words occurring anywhere in the upstream (Re)COGS ”train.tsv”. We also include two words in our vocab/embedding as common nouns accidentally left out of train.tsv vocabulary by the COGS author: ”monastery” and ”gardner” (only included in their train_100.tsv and dev.tsv not also in train.tsv, but present in test/gen), a decision affecting just 22 or 0.1% of generalization examples so would not affect any conclusions qualitatively. See also the discussion on COGS Github with the COGS author at https://github.com/najoungkim/COGS/issues/2#issuecomment-976216841 ., and apply just once per encoder pass 19 attention-head compatible flat pattern-matching rules (Figures 2, 5, 6), shown using grammar coverage (Zeller et al., 2023) to be learnable from the training data [^17] Specific training examples for each rule are in Table 2 the end of ”Appendix: Grammar Coverage analysis to develop and justify Restricted Access Sequence Processing model design” (9.11), plus general pp/cp handling logic. Each pattern handles "det common_noun" and "proper_noun" identically, a symmetry which is evident in the training data. The model does NOT apply recursive, tree-structured rules like ‘np_det pp np -> np_pp -> np‘. Consistent with (Zhou et al., 2023) we use (Weiss et al., 2021)’s RASP originally used for modeling Transformer encoders to model an encoder-decoder in a causal way by feeding the autoregressive output back into the program. [^18] We only have aggregations with non-causal masks when that aggregation (or without loss of generality just before the aggregation product is used to avoid multiplying everywhere) is masked by an input mask restricting it to the sequence corresponding to the input.
For training Transformers from scratch with randomly initialized weights, we use scripts derived from those provided by (Wu et al., 2024)191919 https://github.com/frankaging/ReCOGS/blob/ 1b6eca8ff4dca5fd2fb284a7d470998af5083beb/run_cogs.py and https://github.com/frankaging/ReCOGS/blob/ 1b6eca8ff4dca5fd2fb284a7d470998af5083beb /model/encoder_decoder_hf.py . See "Appendix: Model Detail" (9.4).
5 Methods
We also measure grammar coverage (Zeller et al., 2023) (more detail in Appendix 9.10) by input examples supported by our RASP model against the full grammar of COGS/ReCOGS input sentences provided in the utilities of the IBM CPG project (Klinger et al., 2024) [^23] https://github.com/IBM/cpg/blob/ c3626b4e03bfc681be 2c2a5b23da0b48abe6f570 /src/model/ cogs_data.py#L523 . See "Appendix: Methods Detail" (9.5). See also "Appendix: Results Notebook links by section" (9.1) for notebooks documenting results and giving steps to reproduce.
6 Results
Restricted Access Sequence Processing - grammar coverage using a flat pattern matching approach (not tree-based and not recursive) and autoregressive decoder loop
Figures 2, 5, and 6. See "Appendix: Grammar Coverage Analysis for Design of Restricted Access Sequence Processing Model" (9.11) for more details. We generated 21 sentences based on rules present in the training examples which cover 100% of the COGS input grammar [^24] ReCOGS has same input sentences as COGS, only logical form output is different(lexical differences ignored, under the context-free grammar, tree-based assumption which is violated for our non-tree non-recursive model for prepositional phrases, requiring an additional rule to avoid "attraction errors", Figure 7) per (Zeller et al., 2023). Attention-head compatible flat patterns/rules were derived from those examples (detail in Table 2):
# 19 flat patterns for non-recursive grammar rules
((det common)|proper) was v_trans_omissible_pp_p1
((det common)|proper) v_trans_omissible_p1
((det common)|proper) v_trans_omissible_p2
  ((det common)|proper)
((det common)|proper) was v_trans_omissible_pp_p2
  by ((det common)|proper)
((det common)|proper) v_trans_not_omissible
  ((det common)|proper)
((det common)|proper) was v_trans_not_omissible_pp_p1
((det common)|proper) was v_trans_not_omissible_pp_p2
  by ((det common)|proper)
((det common)|proper) v_unacc_p1
  ((det common)|proper)
((det common)|proper) was v_unacc_pp_p1
((det common)|proper) was v_unacc_pp_p2
  by ((det common)|proper)
((det common)|proper) v_inf_taking to v_inf
((det common)|proper) v_unerg
((det common)|proper) v_unacc_p2
((det common)|proper) v_dat_p1
  ((det common)|proper) to ((det common)|proper)
((det common)|proper) v_dat_p2
  ((det common)|proper) ((det common)|proper)
((det common)|proper) was v_dat_pp_p3
  ((det common)|proper)
((det common)|proper) was v_dat_pp_p4
  ((det common)|proper)
  by ((det common)|proper)
((det common)|proper) was v_dat_pp_p2
  to ((det common)|proper) by ((det common)|proper)
((det common)|proper) was v_dat_pp_p1
  to ((det common)|proper)
# 2 examples for recursive grammar rules
# 1 prepositional phrase example
# (flat rule: mask out "pp ((det common)|proper)"
#             except when outputting noun and nmod)
((det common)|proper) v_trans_omissible_p2
  ((det common)|proper) pp ((det common)|proper)
# 1 sentential complement example
# (flat rule: mask out cp prefix except
#  when outputting that part of LF)
((det common)|proper) v_cp_taking that
  ((det common)|proper)
  v_trans_omissible_p2 ((det common)|proper)
The first 19 of those patterns are present in our RASP program code [^25] https://github.com/willy-b/learning-rasp/blob/ 2be461ba83b7393fda4dc5a96ae600ac98478771 /word-level-pos-tokens-recogs-style-decoder-loop.rasp#L574 and each correspond to a set of RASP operations corresponding to attention operations in a Transformer to match a template corresponding to that sentence type262626see also ”Appendix: Restricted Access Sequence Processing word-level token program/model design” (9.2). To handle prepositional phrases in a flat solution, we find it necessary on the training data to add a rule that ignores "det common_noun" or "proper noun" preceded by a preposition when searching for noun indexes to report in relationships (agent, theme, recipient, etc) and as if we did that during pattern matching by using before/after matches instead of strict relative indexing. Considering how a model without this rule would behave led us to be able to predict 96% of a certain category of errors a baseline Encoder-Decoder Transformer makes (see baseline attraction error results page 1).
Figure 2: Example RASP model flat grammar pattern matching, for np v_dat_p2 np np, for a matching sentence. See Figure 5 and 6 in the Appendix for matching a sentence with middle-noun pp modification and non-matching cases.
ReCOGS_pos Split Semantic Exact Match % (95% CI)
ReCOGS_pos test set (held out, in-distribution) 100.00% (99.88-100.00%)
Generalization splits (held out, out-of-distribution) (below)
active_to_passive 100.00% (99.63-100.00%)
do_dative_to_pp_dative 100.00% (99.63-100.00%)
obj_omitted_transitive_to_transitive 100.00% (99.63-100.00%)
obj_pp_to_subj_pp 92.20% (90.36-93.79%)
obj_to_subj_common 100.00% (99.63-100.00%)
obj_to_subj_proper 100.00% (99.63-100.00%)
only_seen_as_transitive_subj_as_unacc_subj 100.00% (99.63-100.00%)
only_seen_as_unacc_subj_as_obj_omitted_transitive_subj 100.00% (99.63-100.00%)
only_seen_as_unacc_subj_as_unerg_subj 100.00% (99.63-100.00%)
passive_to_active 100.00% (99.63-100.00%)
pp_dative_to_do_dative 100.00% (99.63-100.00%)
prim_to_inf_arg 100.00% (99.63-100.00%)
prim_to_obj_common 100.00% (99.63-100.00%)
prim_to_obj_proper 100.00% (99.63-100.00%)
prim_to_subj_common 100.00% (99.63-100.00%)
prim_to_subj_proper 100.00% (99.63-100.00%)
subj_to_obj_common 100.00% (99.63-100.00%)
subj_to_obj_proper 100.00% (99.63-100.00%)
unacc_to_transitive 100.00% (99.63-100.00%)
all gen splits (19K examples, aggregate) 99.59% (99.49-99.68%)
Table 1: ReCOGS_pos test set performance (n=3000) and non-recursive out-of-distribution generalization split performance for Restricted Access Sequence Processing (RASP) Encoder-Decoder Transformer-compatible model (n=1000 per gen split).
Restricted Access Sequence Processing - test set and generalization set performance
The Restricted Access Sequence Processing program scored 100% Semantic Exact Match and String Exact Match (no missed examples) (95% confidence interval (Beta dist / Clopper-Pearson) of 99.88% to 100%, n=3000) on the ReCOGS_pos test set. The RASP model scored 99.59% semantic exact match on all non-recursive out-of-distribution generalization splits (18922 out of 19000 (95% confidence interval: 99.49% to 99.68%)). See Table 1. Recursion splits are reported below.
Restricted Access Sequence Processing - prepositional phrase and sentential complement recursion (tail recursive) with a non-tree, non-recursive approach using the decoder loop [^27] The grammar includes two (tail) recursive aspects, prepositional phrase and sentential complement recursion. The prepositional phrase recursion comes from the following COGS input grammar rules: ‘np -¿ np_det — np_prop — np_pp‘ and ‘np_pp -¿ np_det pp np‘. Thus np can be expanded in an unbounded way as follows: ‘np -¿ (np_det pp np) -¿ np_det pp (np_det pp np) -¿ np_det pp np_det pp (np_det pp np)‘ and so on. However, one sees this is tail recursion and can be handled by a loop that just appends ‘np_det pp‘ until the final ‘np‘ is not ‘np_pp‘. sentential complement recursion arises from ‘np v_cp_taking that start‘, which can recursively expand as ‘np v_cp_taking that start -¿ np v_cp_taking that (np v_cp_taking that start)‘, and so on until the nonterminal start expands to some other non sentential complement related nonterminal.
Our RASP model’s ReCOGS pp_recursion AND cp_recursion gen split scores were both 100% semantic exact match AND string exact match (95% confidence interval (Beta dist/Clopper-Pearson): 99.63% to 100.0%; n=1000 for each). See Table 1.
(Wu et al., 2024) Encoder-Decoder Transformer from scratch baselines (ReCOGS_pos)
(Wu et al., 2024)’s baseline Encoder-Decoder Transformer on ReCOGS_pos had an overall score of 88.55% +/- 1.87% Semantic Exact Match accuracy (sample +/- std, n=20) with a 95% confidence interval for the sample mean when n=20 of 87.73% to 89.37%.
(Wu et al., 2024)’s baseline Encoder-Decoder Transformer’s Semantic Exact Match score on the extremely difficult obj_pp_to_subj_pp split of ReCOGS_pos was 19.7% +/- 6.1% Semantic Exact Match accuracy (sample +/- std, n=20) with 95% confidence interval for the sample mean with n=20 of 17.0% to 22.4% .
(Wu et al., 2024)’s baseline Encoder-Decoder Transformer’s Semantic Exact Match score on the pp_recursion split of ReCOGS_pos was 40.18% +/- 2.07% Semantic Exact Match accuracy (sample +/- std, n=20) with 95% confidence interval for the sample mean with n=20 of 36.13 to 44.24%. (Wu et al., 2024)’s baseline Encoder-Decoder Transformer’s Semantic Exact Match score on the cp_recursion split of ReCOGS_pos was 52.40% +/- 1.38% Semantic Exact Match accuracy (sample +/- std, n=20) with 95% confidence interval for the sample mean with n=20 of 51.80 to 53.01%.
(Wu et al., 2024) Encoder-Decoder baseline 2-layer Transformer does not improve on the obj_pp_to_subj_pp split when adding 1 or 2 additional layers (even allowing parameter count to increase) [^28] Since no improvement was observed, we did not run the costly experiments to increase the layers while controlling the parameter count (which would be a follow up to distinguish if the improvement was from the layer increase or the parameter increase).
3-layer (Wu et al., 2024) Encoder-Decoder on ReCOGS_pos obj_pp_to_subj_pp split: 16.2% +/- 2.7% Semantic Exact Match (sample mean +/- std, n=10) with 95% confidence interval for sample mean (n=10) of 14.6% to 17.9% . 4-layer (Wu et al., 2024) Encoder-Decoder on ReCOGS_pos obj_pp_to_subj_pp split: 19.3% +/- 4.1% Semantic Exact Match (sample mean +/- std, n=10) with 95% confidence interval for sample mean (n=10) of 16.8% to 21.8%.
Attraction Error Analysis for (Wu et al., 2024) baseline Encoder-Decoder Transformer on obj_pp_to_subj_pp split
(For additional methods detail see Appendix (9.8).) Of the obj_pp_to_subj_pp split single part errors in single verb sentences made by the (Wu et al., 2024) baseline Encoder-Decoder Transformer where the agent was to the left of the verb [^29] Our hypothesis is in terms of nouns with a logical form relationship to a verb or other noun, where the relationship could be agent, theme, or recipient. Since the obj_pp_to_subj_pp split is in terms of subject vs object prepositional modification (instead of agent, recipient, or theme), we use the subset of sentences within this split where the agent is to the left of the verb and pp-modified as it corresponds to the subject in that case. , across n=10 models [^30] On a per model basis (n=10), the fraction of agent-left single point errors where it was the agent relationship in the logical form that was broken were: [0.985, 1.0, 1.0, 1.0, 1.0, 0.990, 1.0, 1.0, 1.0, 1.0]. , 765 out of 767 (99.74%; 95% confidence interval 99.06 to 99.97%) were in the agent part of the logical form (the predicted position for the error).
Critically across all n=10 (Wu et al., 2024) models, for 96.73% (740 out of the previously mentioned 765 above; 95% confidence interval (Beta dist / Clopper-Pearson) 95.21 to 97.87%) of the single point errors in logical forms for single verb sentences where the agent was on the left, modified by a prepositional phrase, and the error was in the agent part, the error in the logical form was that the agent index was accidentally assigned to the specific expected prepositional phrase noun (the one closest to the verb on the left side) instead of the original agent noun. (Figure 7) This does not vary much from randomly initialized and trained model to model, with the model-level average at 97.07% of such errors exactly as predicted (stderr=2.23% (n=10)), with 7 of 10 models having 100% of these errors exactly as predicted by our hypothesis [^31] Fraction for each model as predicted: [0.970, 0.761, 1.0, 1.0, 1.0, 1.0, 1.0, 1.0, 0.976, 1.0]). . The attraction to the nearest noun hypothesis predicts that the offset in the agent index varies with prepositional phrase recursion depth (as at depth > 1, there are multiple attractor prepositional nouns to choose from).  [^32] See ”Appendix: Attraction errors” (9.6) for examples of different pp recursion depths.
We report that for all (n=22) single logical form part errors observed (from running n=10 separate Transformer models over the 1000 sentences in the split) where in the input the agent is left of the verb and has a depth=2 prepositional phrase modification in this split, in 100% (95% confidence interval (Beta dist / Clopper-Pearson) 84.6 to 100%; n=22) of those sentences the error in the agent right-index matched our prediction.
(Wu et al., 2024) Encoder-Decoder Transformer on new v_dat_p2 pp moved to recipient (from theme) split - as hard as hardest previous generalization split
See Figure 3 . As the RASP model predicted the ‘np v_dat_p2 np pp np np‘ [^33] Strictly speaking we only do ‘np v_dat_p2 np_det pp np np‘ as per the grammar ‘np_prop‘ cannot precede a prepositional phrase prepositional phrase modification (which involves the recipient instead of the subject so is a distinct check of our hypothesis) would require learning to ignore the distractor "pp det common_noun" and "pp proper_noun" same as required for the obj_pp_to_subj_pp split, we predicted that a new split we introduce "v_dat_p2_pp_moved_to_recipient" would also be difficult for the Transformer. To test this, (Wu et al., 2024)’s baseline Encoder-Decoder Transformer was trained with default data (ReCOGS_pos train.tsv) and tested on modified v_dat_p2 pp training examples where only the word order was changed to move the prepositional phrase from the theme to the recipient (logical form properly updated see Appendix 9.9 for all examples). The baseline (Wu et al., 2024) Encoder-Decoder Transformer was only able to achieve a Semantic Exact Match (sample mean +/- sample std) of 13% +/- 15.6% (n=10 Transformers trained from scratch with randomly initialized weights and data shuffling) with a 95% confidence interval for the sample mean when n=10 of 4% to 23%. Thus, this new split we introduce here as v_dat_p2_pp_moved_to_recipient is as difficult or perhaps more difficult than the previous reported "hardest split" obj_pp_to_subj_pp.
(Wu et al., 2024) Encoder-Decoder Transformer trained with data augmented with v_dat_p2 pp moved to recipient (from theme) does NOT improve obj_pp_to_subj_pp performance
(Wu et al., 2024)’s baseline Encoder-Decoder Transformer was trained with default data (ReCOGS_pos train.tsv) but with additionally the same modified v_dat_p2 pp training examples used for the "v_dat_p2_pp_moved_to_recipient" split (non-subject recipient modified with prepositional phrase, so nonoverlapping with subj_pp, Figure 3 (b) and (d)) above on which it performed poorly, then tested on the standard prepositional modification generalization split "obj_pp_to_subj_pp", after which it achieved 22% +/- 6.7% Semantic Exact Match (sample mean +/- std, n=10) with 95% confidence interval for sample mean n=10 of 17.9% to 26.1% (not significantly different than (Wu et al., 2024)’s baseline by one-tailed Welch’s unequal variances t-test).
7 Analysis
Our RASP model of a Transformer Encoder Decoder, without tree-based or recursive aspects scored 100% in semantic exact match accuracy on the (Wu et al., 2024) test set (n=3000), and on the generalization data scored 100% in all but one category (see above) without explicit rules in the RASP program to handle them. This includes 100% semantic exact match accuracy on the prepositional phrase recursion and sentential complement recursion generalization splits up to depth 12 (n=1000 examples each), without any hardcoded prepositional phrase or sentential complement expansion shortcuts added343434a single rule applies to all depths; the only limit on length generalization is the RASP interpreter and a simple to extend positional encoding which only handles sentences up to a limit . The RASP program only made a significant number of errors on obj_pp_to_subj_pp which scored only 92.20% Semantic Exact Match (95% confidence interval (Beta dist / Clopper-Pearson): 90.36% to 93.79%) Semantic Exact Match accuracy, much better than (Wu et al., 2024) baseline Encoder-Decoder Transformers which only scored 19.7% +/- 6.1% Semantic Exact Match (sample mean +/- std) with 95% confidence interval for the sample mean with n=20 of 17.0% to 22.4% (n=20 separately trained models with different random seeds for weight initialization and training data ordering; n=1000 examples used to test each of the n=20 models). Thus, we demonstrated by construction using the Restricted Access Sequence Processing language which can be compiled to concrete Transformer weights that theoretically a Transformer Encoder-Decoder can solve the COGS input to ReCOGS_pos logical form translation in a systematic, compositional, and length generalizing way.
Recall we found a single flat pattern matching rule we originally added to fit training examples, to ignore "pp det common_noun" and "pp proper_noun" when matching nouns for the agent, theme, recipient right indices, was sufficient to avoid structural generalization errors due to pp modification in novel positions.
Interestingly, we imagined ablating that single rule and hypothesized attraction to the nearest noun (Figure 7 and "Appendix: Attraction Errors" (9.6)) in its absence and found this predicted the exact error (the nearest noun to the verb on the expected side is mistaken for the agent of the verb) in 96% of the single relationship errors the (Wu et al., 2024) baseline Transformers make on the obj-pp-to-subj-pp split353535overall, their semantic exact match on the split is measured by us at 19.7%, consistent with their Figure 5 when the agent is left of the verb in single verb sentences (suggesting perhaps that the baseline (Wu et al., 2024) Transformer trained from scratch is also not learning a hierarchical, tree-structured representation.) [^36] We found (Li et al., 2023) also observe this stating ”For instance, in sentences like ’A cat on the mat froze’, models often misinterpret the closer NP the mat as the subject.” Our explanation could have been refuted by other single relationship errors occurring as frequently as the agent, indicating general model confusion (independently getting incorrect agent and theme, not just agent relationships) and/or when making an agent error, the model could have simply put nonsense indices or referred to any other position other than the closest pp noun position to the verb (which does vary and depends on pp depth) to refute our hypothesis.
The flat/non-tree structured hypothesis for the baseline Transformer can also be checked by making a prediction on a completely different syntax affected by the same issue: the ‘np v_dat_p2 np pp np np‘ prepositional phrase modification (which involves the recipient relationship being modified instead of the subject and/or agent so is a distinct check of our hypothesis) [^37] See Appendix 9.9 and we indeed found that this was as hard or harder than the previous most difficult split analyzed above, the ‘obj_pp_to_subj_pp‘ split (Figure 3).
Figure 3: (Wu et al., 2024) Encoder-Decoder Transformer trained from scratch generalizing to new v_dat_p2 pp moved to recipient (from theme) split is as hard as the previously reported hardest generalization split consistent with the flat/non-recursive/non-tree representation hypothesis ((d) rejects $H_{0}$). Figure 5 shows how a flat RASP model can recognize (b).
Maybe (Wu et al., 2024) baseline Encoder-Decoder is depth-constrained to learn a non-hierarchical, flat, non-tree model with these characteristic errors and with more layers it would learn to recursively combine ‘np_det pp np -> np_pp -> np‘ (to some fixed depth at least, probably limited by the number of Transformer blocks) and perform better on prepositional phrase related splits383838(Csordás et al., 2022): ”the network should be sufficiently deep, at least as deep as the deepest data dependency in the computational graph built from elementary operations (e.g., in the case of a parse tree, this is the depth of the tree)”. (This is not a very scalable approach, as we must make the network deeper to handle increased pp depths instead of just looping or at least reusing weights each layer.). However, training a (Wu et al., 2024) baseline Encoder-Decoder Transformer from scratch we found no benefit to 3 or 4 layers instead of 2 on the ReCOGS obj_pp_to_subj_pp split, consistent with (Petty et al., 2024)’s finding on COGS.
Taken together, these results and the grammar coverage analysis suggest we may interpret the poor performance on generalizing on unseen prepositional phrase related modification related splits as arising from the baseline 2 to 4 layer Encoder-Decoder Transformers learning a non-tree, non-recursive representation that cannot leverage the grammar rule ‘np_det pp np -> np_pp -> np‘ during learning and which requires them to instead actually observe more of the various prepositional phrase substitutions to learn them. [^39] See ”Appendix: Composition and Learning” (9.13)
8 Conclusion
Implementing our task in Restricted Access Sequence Processing immediately helped us discover additional failure modes (e.g. new "v_dat_p2_pp_moved_to_recipient" split [^40] After this paper was written we found our predicted split ”v_dat_p2_pp_moved_to_recipient” has also been added to an extended separate (SLOG) version of COGS (upstream of ReCOGS) recently in (Li et al., 2023) see their section 2.2.1 indirect object modification (4), confirmed difficult) of (Wu et al., 2024)’s baseline Encoder-Decoder Transformer, predict errors in detail made in the logical forms, and demonstrated that tree-structured/recursive representations are not needed to cover the recursive aspects of the COGS/ReCOGS grammar414141same input grammar is used by COGS and ReCOGS. We recommend others to consider RASP to understand Transformer behavior even for more complicated tasks like ReCOGS. We predict that Transformers should be able to perform the ReCOGS task (even structural generalization splits) with high accuracy and that the problem is just of getting the Transformer to learn the appropriate rules424242e.g. to ignore ”pp det common_noun” and ”pp proper_noun” when finding nouns in relationships with verbs, which allows the RASP model to get 100% on the ReCOGS test set, and 100% two of the three structural generalizations, and 92% on the obj-pp-to-subj-pp split, turning attention to data augmentation [^43] We tried one augmentation, see Results and Appendix 9.9, but there are many other possibilities. , curriculum learning (Bengio et al., 2009), reinforcement learning (Ranzato et al., 2016), training objectives (Ahuja et al., 2024), and other approaches (Csordás et al., 2021) (Ontanon et al., 2022).
Known Project Limitations
The Restricted Access Sequence Processing code is not optimized. Cannot yet predict attention heads and layers required from the select and aggregate operations performed like the RASP authors (Weiss et al., 2021) were able to do with their problems.
Grammar coverage (Zeller et al., 2023) is only valid when the expansions are rules your model can learn. [^44] If for example, as with our flat RASP model by design or as we hypothesize for (Wu et al., 2024)’s baseline Encoder-Decoder Transformer, the model cannot or will not learn the rule ‘np_det pp np -¿ np_pp -¿ np‘ which recursively replaces noun phrases modified by prepositional phrases with a noun phrase, then grammar coverage will still assume any prepositional phrase exposure is sufficient, which is evidently not true given the errors on prepositional phrase modification generalization splits reported here and by (Wu et al., 2024), (Kim and Linzen, 2020). We specifically made use of this limitation in this paper but still caution anyone about it who might just take the grammar coverage metric away by itself.
The attraction error analysis of the (Wu et al., 2024) baseline Encoder-Decoder Transformer on the obj_pp_to_subj_pp split does not yet attempt to explain the common case of multiple errors in the logical form.454545e.g. agent index may be replaced by prepositional phrase noun but also a spurious theme relationship is added or the theme index is also corrupted
We only provide and discuss a RASP solution for ReCOGS (Wu et al., 2024) here, not the semantically equivalent COGS464646nor the recently introduced extended version SLOG (Li et al., 2023) (Kim and Linzen, 2020), though as this goes to publication we have just separately released a RASP model for COGS at https://github.com/willy-b/RASP-for-COGS/ [^47] RASP-for-COGS also supports case-sensitive string exact match and ignoring out-of-vocabulary words, features not supported in RASP-for-ReCOGS, which is undergoing evaluation with preliminary data supporting the same conclusions for that task (can also be solved by a non-tree structured, non-hierarchical Transformer compatible model, despite using Exact Match instead of Semantic Exact Match, with same RASP for the Encoder as used here for ReCOGS_pos but a different Decoder).
Much deeper Transformer networks may be learning a tree-based grammar representation [^48] Nothing explored here rules that out and there is plenty of evidence outside the COGS task-related literature suggesting this will be the case: (Tenney et al., 2019) show the 24-layer BERT model seems to handle ”POS tagging, parsing, NER, semantic roles, then coreference”; (Hewitt and Manning, 2019) ”provid[e] evidence that entire syntax trees are embedded implicitly in deep models’ [including BERT’s] vector geometry”, and (Goldberg, 2019) shows BERT excels at subject-verb agreement, ”which [is] traditionally taken as evidence for the existence [of] hierarchical structure” (though e.g. in this work we see that ignoring distractor nouns in long-term dependencies does not require hierarchy or a deep understanding of syntax but simple rules like ignore ”pp det common_noun” and ”pp proper_noun” for finding noun-verb relationships in the logical form can allow for handling of such long-range dependencies). On the other hand, (Petty et al., 2024) argue specifically for the COGS benchmark (semantically equivalent to ReCOGS which is derived from it) that increasing depth does not allow their Transformers to make progress on the structural generalization splits, even at depths up to 32 layers. and not suffer from the predicted generalization issues observed in (Wu et al., 2024)’s baseline 2-layer Transformer and predicted by our intentionally non-tree RASP model (if compensating rules to avoid attraction errors in a flat model are not also learned).
As stated elsewhere in the paper, we are focused on the structural generalizations that Transformers currently struggle with (prepositional phrase recursion (pp recursion), sentential complement (cp) recursion, object-pp to subj-pp modification generalization (obj-pp-to-subj-pp)) and not lexical generalizations (which Transformers are demonstrated to do well on by others). Thus, as noted elsewhere, our model assumes the embedding can map words to their possible part-of-speech and verb-type and does not address the learnability of this mapping (number of exposures required, etc).
The ReCOGS task is English specific and our findings on compositional generalization may not necessarily apply (or apply differently) to other languages. Our description of "attraction errors" for example does not appear specific to subject-verb-object ordering in English but a similar analysis should be conducted in other languages and we used the obj-pp-to-subj-pp split in our detailed error analysis which is subject-verb-object ordering sensitive.
The RASP-for-ReCOGS model applied to unintended use WILL give invalid results or halt - we have NOT provided a general language model, we have provided a simulation of how a Transformer could perform a specific task. The RASP-for-ReCOGS model/simulation as provided is for research purposes only to prove feasibility of the ReCOGS task by Transformers and is not appropriate for ANY other uses whatsoever without modification. For one, an actual Transformer performing the equivalent operations would run orders of magnitude faster, which should be reason enough to not want to use the RASP simulation for actual input-output tasks outside of a research setting. However, there is also no "tokenizer" provided for the RASP model to handle out-of-vocabulary inputs and fallback paths for out-of-grammar examples are not provided so the RASP model will halt on most inputs and can only run on the in-distribution (non-augmented) training data, and the dev, test, and gen sets of ReCOGS, though such aspects could be added. We provide the code for reproducing the results of this study and for researchers who are capable of writing RASP themselves to build upon the work and/or more easily apply RASP to their own problems given our examples, not for immediate application to any other tasks without appropriate modification.
See also "Appendix: Potential Risks" (9.14) and "Appendix: Scientific Artifacts - Is Our Use Consistent with Authors’ Intention and Licensing" (9.15) for other potential concerns.