### v3

# Exploring Compositional Generalization (in COGS/ReCOGS_pos) by Transformers using Restricted Access Sequence Processing (RASP)

Abstract
Humans understand new combinations of words encountered if they are combinations of words recognized from different contexts, an ability called Compositional Generalization. The COGS benchmark (Kim and Linzen, 2020) reports 0% accuracy for Transformer models on some structural generalizations. We use (Weiss et al., 2021)’s Restricted Access Sequence Processing (RASP), a Transformer-equivalent programming language, to demonstrate that a Transformer Encoder-Decoder can perform COGS and the semantically equivalent ReCOGS_pos(Wu et al., 2024) systematically and compositionally: Our RASP models attain near perfect scores on structural generalization splits on COGS (exact match) and ReCOGS_pos (semantic exact match). Our RASP models show the (Re)COGS tasks do not require a hierarchical or tree-structured solution (contrary to (Kim and Linzen, 2020), (Yao and Koller, 2022), (Murty et al., 2022), (Liu et al., 2021)): we use word-level tokens with an "embedding" layer that tags with possible part of speech111per (Tenney et al., 2019) by layer 0 the part-of-speech could be predicted for most words in Transformers trained on a masked language modeling objective, so we assume an equivalent embedding is learnable., applying just once per encoder pass 19 attention-head compatible flat pattern-matching rules (easily identified with specific training examples), shown using grammar coverage (Zeller et al., 2023) to cover the non-recursive aspects of the input grammar, plus masking out prepositional phrases (’pp noun’) and/or sentential complements (cp) when recognizing grammar patterns and extracting nouns related to the main verb in the sentence, and output the next logical form (LF) token (repeating until the LF is complete). The models do not apply recursive, tree-structured rules like ‘np_det pp np -> np_pp -> np‘, but score near perfect semantic and string exact match on both COGS and ReCOGS pp recursion, cp recursion using the decoder loop.
1 Introduction
It was long argued that connectionist models (i.e. neural networks) were incapable of compositional generalization (Fodor and Pylyshyn, 1988). [^2] More specific versions of this debate continue, for example re: syntax, one can read (van Schijndel et al., 2019) vs (Goldberg, 2019) or re: hierarchical generalization by Transformers, (Petty and Frank, 2021) vs (Murty et al., 2023). However, large language models based on the Transformer architecture (Vaswani et al., 2017) compose seemingly fluent and novel text and are excellent few or zero shot learners (Brown et al., 2020).
Some observations do contradict that Transformers learn systematic, compositional solutions to problems that generalize, for example structural generalization performance on the COGS task (Kim and Linzen, 2020) and ReCOGS variant (Wu et al., 2024) , benchmarks based on extracting semantics (logical form) from the syntax (grammatical form) of synthetic sentences in a simplified subset of English grammar, requiring models trained only on certain grammar examples to generalize to sentences with unseen grammar built up / recombined from parts present in the training examples.
We use (Weiss et al., 2021)’s Restricted Access Sequence Processing (RASP) language that can be compiled to concrete Transformer weights to prove by construction that a Transformer Encoder-Decoder [^3] We follow (Zhou et al., 2023) and (Friedman et al., 2025) who used RASP to analyze auto-regressive decoder-loop cases, not just Transformer encoders as in (Weiss et al., 2021). (hereafter EncDec) can perform COGS and ReCOGS_pos444official variant, closer to COGS than non-positional ReCOGS, as COGS is also positional, and means we can also measure string exact match, not just semantic exact match over the vocabulary and grammar of that task in a systematic, compositional way as a rigorous starting point to investigating when Transformers might learn or not actually learn such compositional/systematic solutions. We find a flat, not hierarchical/tree-structured model which lacks any handling for the recursive rules in the grammar (for prepositional phrase recursion and sentential complement recursion) can perform the task accurately, but requires learning a masking rule555to mask out prepositional phrases when extracting noun-verb relationships (which are separable from pp relationship output in the (Re)COGS logical forms), see center-embedded pp examples in the (Re)COGS training set, see Table 1. for avoiding "attraction" errors [^6] These attraction errors are similar to those discussed elsewhere in NLP and psycholinguistics literature on hierarchical vs linear processing by language models and humans, see ”Appendix: Attraction errors” (9.6). where inserted prepositional phrase nouns replace agent/theme/recipient nouns in the logical form by accident. This is our main result and suggests that Transformers can perform the (Re)COGS task accurately even for novel prepositional phrase substitution positions or structural recursion depths unseen in training, turning efforts to learnability of such a generalizing solution (flat pattern matching with masking), and also adds to the literature a caveat on interpreting success on (Re)COGS by noting a hierarchical or tree-structured representation is not necessarily required (contrary to (Kim and Linzen, 2020) and assumption of (Murty et al., 2022)). Finally, we find that these "attraction" errors predicted for flat pattern matching without masking (and not for recursive/hierarchical/tree-structured representations) are contributing to the high error rate of the (Wu et al., 2024) baseline Transformer trained from scratch (EncDec only 2 layers deep).
2 Prior Literature
(Kim and Linzen, 2020) introduce the COmpositional Generalization Challenge based on Semantic Interpretation (COGS) benchmark, arguing that Transformers have low accuracy on the generalization splits (35% overall), especially structural generalization splits where near 0% accuracy is reported, using a 2-layer EncDec Transformer (2 layers for Encoder, 2 layers for Decoder), and speculate that a hierarchical/tree-structured inductive bias is likely to be required to overcome the difficulty in structural generalization. (Murty et al., 2022) study how Transformers learn to solve COGS and other compositional datasets and measure the tree-structuredness of the representations and argue it predicts which training checkpoints exhibit better compositional generalization. (Yao and Koller, 2022) survey seq2seq models broadly (both LSTMs and Transformers) and find they all struggle on COGS structural generalizations and give two examples of explicitly tree-structured models (forcing tree-structured representations) as an alternative that do very well (Liu et al., 2021)777(Liu et al., 2021) uses two modules, one that ”learns to model the latent syntactic algebra [using a Tree-LSTM to] produce the latent syntactic structure of each expression” and a second that ”learns to assign semantic operations to syntactic operations” so that they can ”transform a syntactic tree to the final composed semantic meaning”, and in this way achieves 97.7% on COGS. However this is neither a Transformer architecture nor a vanilla unstructured neural network, it uses an LSTM specifically architected based on the assumption of a recursive, tree-structured context-free grammar. Note LeAR also uses what appear to be heavily crafted semantic primitives in their Interpreter module, see their Table 8 for COGS, requiring manual task specific design. and (Weißenhorn et al., 2022). Finally, (Lake and Baroni, 2023) use a "meta-learning for compositionality" approach with a 3-layer EncDec Transformer architecture and achieve what they call "human-like systematic generalization", achieving high scores on everything in the COGS benchmark (>99% on lexical generalizations) EXCEPT the structural generalization splits where they also still score 0% accuracy. However, one notices these networks are shallow compared with those used in successful large-pretrained Transformer models (e.g. 24-layer BERT where compositional parse trees seem to be encoded in its vector space representation (Hewitt and Manning, 2019)), and it is claimed, by e.g. (Csordás et al., 2022) that for compositional operations, like parsing, the depth of the network must be at least the maximum number of compositional operations, e.g. the height of the parse tree. Remarkably, (Petty et al., 2024) finds that increasing the layer depth of the Transformer models (up to 32 layers) does not improve the near 0% accuracy on COGS structural generalization splits like prepositional phrase modification of subject when the network has only seen it on the object during training and also input length/depth generalizations (like pp/cp recursion), perhaps surprising as for the simpler logical inferences problem in (Clark et al., 2020) they observed successful logical inference depth generalization even by Transformer Encoder-only Transformers.
Thankfully, (Wu et al., 2024) are able to begin to get traction (low but nonzero accuracy) for the shallow EncDec Transformer models on structural generalizations in a modified but semantically equivalent form of the COGS task they call ReCOGS. They remove redundant symbols, and use Semantic Exact Match instead of Exact Match (Fig 1).
(Zhou et al., 2023) apply (Weiss et al., 2021)’s RASP language to explain some inconsistent findings regarding generalization, using RASP to predict some cases of generalization that come easily to Transformers and some which do not.
(Zhou et al., 2023) seem to reveal (Weiss et al., 2021) has provided the framework we seek by demonstrating how to apply RASP to Transformer decoders with intermediate steps, and even use it to learn how to modify difficult-to-learn tasks like Parity [^8] See (Strobl et al., 2024) for context from formal language theory, computational complexity, circuit complexity theory, and experimental papers together, providing robust lower and upper bounds on what Transformers can do, including discussion of under what conditions Parity can be solved by Transformers and how whether it can be learned by randomly initialized Transformers under simple training schemes is a different question (general feed-forward neural networks can learn to solve Parity per (Rumelhart et al., 1988)). (Delétang et al., 2023) also. See Appendix (9.12) and long addition in seemingly incidental ways based on RASP analysis to make them readily learnable by Transformers in a compositional, length generalizing way!
Thus we apply techniques similar to (Zhou et al., 2023) and (Weiss et al., 2021) to (Re)COGS to (1) argue Transformers should be able to perform the (Re)COGS task accurately even for novel prepositional phrase substitution positions or structural recursion depths unseen in training, and that the problem is learning not capability and (2) try to understand the prepositional phrase modification related generalization errors (Wu et al., 2024)’s baseline EncDec Transformers are making.
Figure 1: Introducing parse trees, logical form, and semantic graphs. Two semantically identical but syntactically distinct (i) sentences (a) "a boy painted the girl" and (b) "the girl was painted by a boy" are shown with (ii) their distinct parse tree (parsed into the shared (Re)COGS input grammar), (iii) the string form of their semantics (ReCOGS logical form; differs in indices and ordering), and (c) the graph representation of their logical form (semantic graph [^10] As a convention, in converting ReCOGS logical forms to Semantic Graphs (used for our illustrations and to explain Semantic Exact Match, not in the code) we use the logical form (source, target) index order for directed semantic graph edges (from verb to related entity) EXCEPT for the agent relationship which is from the agent of a verb to the verb (opposite direction from logical form in that case), which gives our semantic graphs of ReCOGS sentences an unambiguous starting point (layout starts from agent) without affecting comparison of the graphs (generated by a consistent rule), see also Figure 3. For more information on equivalence of (Re)COGS LF with semantic graphs see e.g. Section 3.2 of (Weißenhorn et al., 2022) who use a semantic graph based representation in their model (they use different conventions but show such mappings exist)., not different at all between the two examples). Note the (iii) logical forms (LFs) differ by String Exact Match but not (Wu et al., 2024)’s Semantic Exact Match (order and indices do not match but nouns, normalized verbs, and relationships between nouns and verbs are same). Note the "agent", "theme" order in the logical form string is not required to match for Semantic Exact Match. COGS and ReCOGS tasks require extracting the semantics (c) encoded in LFs (iii) of sentences (i). RASP-for-(Re)COGS shows decoding the LF (Figures 7 (ReCOGS), 10 (COGS)) does NOT require representing the parse tree.
Figure 2: RASP-for-(Re)COGS Encoder determines an example contains "np v_dat_p2 np np" via flat pattern matching (maps to agent, recipient, theme noun-verb relation ordering (1)). See Figure 5 for middle-noun $p$ modification, Figure 6 for the non-matching case, and Figure 7 (ReCOGS), 10 (COGS) for the full Encoder-Decoder models translating input to logical form. 1d embed. equiv. to higher dim orthogonal vector idx (A9.2).
3 Data
Figure 1. COGS [^11] https://github.com/najoungkim/COGS (Kim and Linzen, 2020) and ReCOGS [^12] https://github.com/frankaging/ReCOGS (Wu et al., 2024) based on COGS (Kim and Linzen, 2020) (input sentences/grammar are the same but logical form output are different) datasets were used, with special attention on the structural generalization splits (especially prepositional phrase Object-to-Subject generalization). The grammar and vocabulary description for COGS English input sentences provided in the utilities associated with the IBM CPG project (Klinger et al., 2024) [^13] https://github.com/IBM/cpg/blob/ c3626b4e03bfc681be2c2a5b23da0b48abe6f570 /src/model/cogs_data.py#L523 were used in designing our RASP solution and analyzing the ways in which this task could be learned (we did not actually use their grammar though, and our RASP solution is flat and non-hierarchical unlike their description of the COGS probabilistic context-free grammar which is hierarchical and recursive).
4 Model
Figures 2 ((Re)COGS shared encoder), 7 (ReCOGS EncDec) , 10 (COGS EncDec). Model code is available: RASP-for-COGS and RASP-for-ReCOGS [^14] https://github.com/willy-b/RASP-for-COGS (COGS) and https://github.com/willy-b/learning-rasp (ReCOGS). We used the RASP interpreter of (Weiss et al., 2021). For RASP model design and details see Appendix 9.2. We use word-level tokens for all RASP model results in this paper, [^15] We believe any solution at the word-level can be converted to a character-level token solution (see Appendix 9.3). with an "embedding" layer1616161d scalar embedding, the default in RASP. Equivalent to specifying an orthogonal vector by index in a higher dimensional space. E.g. embedding to common_noun (7) is as setting 7th dimension to 1 in one-hot, or just using 7th orthogonal vector (given some numbering so we can assign one to each POS and verb type). Equality checks for the 1d scalars are equivalent to vector dot product on the indexed orthogonal vectors. Checking for EITHER common_noun OR proper_noun can use a dot product with the non-orthogonal vector given by the sum of (common_noun + proper_noun). that tags with possible part of speech [^17] Note we followed the (Klinger et al., 2024) description of COGS and include in our RASP-for-(Re)COGS vocabulary (part-of-speech or verb-type embedding/mapping) all words occurring anywhere in the upstream (Re)COGS ”train.tsv”. For RASP-for-ReCOGS only, we also include two words in our vocab/embedding as common nouns accidentally left out of train.tsv vocabulary by the COGS author (COGS vocab is used by ReCOGS): ”monastery” and ”gardner” (only included in their train_100.tsv and dev.tsv not also in train.tsv, but present in test/gen), a decision affecting just 22 or 0.1% of generalization examples so would not affect any conclusions qualitatively. For RASP-for-COGS which was started after RASP-for-ReCOGS, we considered these words out-of-vocabulary to be more conservative. See also the discussion on COGS Github with the COGS author at https://github.com/najoungkim/COGS/issues/2 #issuecomment-976216841 ., and apply just once per encoder pass 19 attention-head compatible flat pattern-matching rules (Table 1, Figures 2, 5, 6; official training examples the rules can be derived from given in Table 1), shown using grammar coverage (Zeller et al., 2023) to cover the non-recursive aspects of the input grammar, plus masking out all prepositional phrases and/or sentential complements (cp) when extracting nouns related to the main verb in the sentence (derived from center embedded pp in training) and separately unrolling local prepositional phrase noun modifiers (Figure 8) / local sentential complement nouns and verb relationships (Figure 9) one by one in the decoder loop. Each pattern handles "det common_noun" and "proper_noun" identically, a symmetry which is evident in the training data. The model does NOT apply recursive, tree-structured rules like ‘np_det pp np -> np_pp -> np‘.
Consistent with (Zhou et al., 2023) we use (Weiss et al., 2021)’s RASP originally used for modeling Transformer encoders to model an encoder-decoder in a causal way by feeding the autoregressive output back into the program. [^18] We only have aggregations with non-causal masks when that aggregation (or without loss of generality just before the aggregation product is used to avoid multiplying everywhere) is masked by an input mask restricting it to the sequence corresponding to the input.
Official COGS/ReCOGS training sentence POS/verb-type sequence matched Noun-verb relationship template
The donut was studied . ((det common)|proper) was v_trans_omissible_pp_p1 theme
The captain ate . ((det common)|proper) v_trans_omissible_p1 agent
The sailor dusted a boy . ((det common)|proper) v_trans_omissible_p2 ((det common)|proper) agent, theme
A drink was eaten by a child . ((det common)|proper) was v_trans_omissible_pp_p2 by ((det common)|proper) theme, agent
A girl liked the raisin . ((det common)|proper) v_trans_not_omissible ((det common)|proper) agent, theme
The pen was helped . ((det common)|proper) was v_trans_not_omissible_pp_p1 theme
A rose was helped by a dog . ((det common)|proper) was v_trans_not_omissible_pp_p2 by ((det common)|proper) theme, agent
A cat disintegrated a girl . ((det common)|proper) v_unacc_p1 ((det common)|proper) agent, theme
A box was inflated . ((det common)|proper) was v_unacc_pp_p1 theme
The cake was frozen by the giraffe . ((det common)|proper) was v_unacc_pp_p2 by ((det common)|proper) theme, agent
The girl needed to cook . ((det common)|proper) v_inf_taking to v_inf agent
The sailor laughed . ((det common)|proper) v_unerg agent
A cake rolled . ((det common)|proper) v_unacc_p2 theme
A girl was sold the cake . ((det common)|proper) was v_dat_pp_p3 ((det common)|proper) recipient, theme
The girl was lended the balloon by Harper . ((det common)|proper) was v_dat_pp_p4 ((det common)|proper) by ((det common)|proper) recipient, theme, agent
The pen was offered to the girl by Emma . ((det common)|proper) was v_dat_pp_p2 to ((det common)|proper) by ((det common)|proper) theme, recipient, agent
The melon was lended to a girl . ((det common)|proper) was v_dat_pp_p1 to ((det common)|proper) theme, recipient
Liam forwarded the girl the donut . ((det common)|proper) v_dat_p2 ((det common)|proper) ((det common)|proper) agent, recipient, theme
Emma passed a cake to the girl . ((det common)|proper) v_dat_p1 ((det common)|proper) to ((det common)|proper) agent, theme, recipient
Center-embed. PP training example RASP-for-(Re)COGS model $\rightarrow$ mask "PP noun" rule (when finding NV rels.)
Isabella forwarded a box on a tree to Emma . ((det common)|proper) v_dat_p1 ((det common)|proper) _(masked pp det common_noun)_ to ((det common)|proper) agent, theme, (masked), recipient
Table 1: 19 flat part of speech / verb type patterns (matched by Encoder, Figures 2, 7) derived from training examples are sufficient to cover the (Re)COGS input grammar after pp/cp masking (Figures 8, 9) to obtain noun-verb relationship ordering for the main verb.
ReCOGS_pos Split Semantic Exact Match % (95% CI)
ReCOGS_pos test set (held out, in-distribution) 100.00% (99.88-100.00%)
Structural generalization splits (held out, out-of-distribution)
obj_pp_to_subj_pp 92.20% (90.36- 93.79%)
pp_recursion 100.00% (99.63-100.00%)
cp_recursion 100.00% (99.63-100.00%)
all gen splits (21K examples, aggregate) 99.63% (99.54-99.71%)
Table 2: RASP-for-ReCOGS Encoder-Decoder Transformer-compatible model performance on the ReCOGS_pos test set (n=3000) and out-of-distribution structural generalization split performance (n=1000 per gen split). Model is deterministic, so Clopper-Pearson CIs are used. (Model was developed on training data or examples identical to training examples after embedding words to part of speech and verb type sequences, the first step in the model, see Tables 1 , 7). 100.00% (CI 99.63-100.00%) string exact match was also attained for recursion splits. See Table 5 for the rest of the 21 generalization splits (only 3 structural shown).
ReCOGS_pos Split Semantic Exact Match % +/- sample std 95% CI (of sample mean)
Structural generalization splits (held out, o.o.d.)
pp_recursion (up to depth 12 vs 0-2 in training) 40.2% +/- 9.3% 36.1-44.2%
cp_recursion (up to depth 12 vs 0-2 in training) 52.4% +/- 1.4% 51.8-53.0%
obj_pp_to_subj_pp (N=20) 19.7% +/- 6.1% 17.0-22.4%
$\rightarrow$ + 1 layer (N=10) 16.2% +/- 2.7% 14.6-17.9%
$\rightarrow$ + 2 layers (N=10) 19.3% +/- 4.1% 16.8-21.8%
Table 3: Out-of-distribution (o.o.d.) ReCOGS_pos structural generalization split performance (n=1000 per gen split) for N=20 (Wu et al., 2024) baseline Encoder-Decoder Transformer models trained from scratch on random perm. of the training set and random weight init. N=10 for +layer expts only. Results consistent with (Wu et al., 2024)’s Fig 5.
COGS Split String Exact Match % (95% CI)
COGS test set (held out, in-distribution) 99.97% (99.81-99.99%)
Structural generalization splits (held out, out-of-distribution)
obj_pp_to_subj_pp 100.00% (99.63-100.00%)
pp_recursion 98.40% (97.41-99.08%)
cp_recursion 99.90% (99.44-99.997%)
all gen splits (21K examples, aggregate) 99.89% (99.83-99.93%)
Table 4: RASP-for-COGS Encoder-Decoder Transformer-compatible model performance on the COGS test set (n=3000) and out-of-distribution structural generalization split performance (n=1000 per gen split). Model is deterministic, so Clopper-Pearson CIs are used. See Table 6 for the rest of the 21 generalization splits (only 3 structural shown).
For training Transformers from scratch with randomly initialized weights, we use scripts derived from those provided by (Wu et al., 2024) [^19] https://github.com/frankaging/ReCOGS/blob/ 1b6eca8ff4dca5fd2fb284a7d470998af5083beb/run_cogs.py and https://github.com/frankaging/ReCOGS/blob/ 1b6eca8ff4dca5fd2fb284a7d470998af5083beb /model/encoder_decoder_hf.py . See "Appendix: Model Detail" (9.4).
5 Methods
We use the RASP (Weiss et al., 2021) interpreter202020provided at https://github.com/tech-srl/RASP/ to evaluate our RASP programs. Logical forms (LFs) generated by the models were scored by String Exact Match (COGS) and Semantic Exact Match [^21] Used official scripts https://github.com/frankaging/ReCOGS/blob/ 1b6eca8ff4dca5fd2fb284a7d470998af5083beb/utils/ train_utils.py and https://github.com/frankaging/ReCOGS/blob/ 1b6eca8ff4dca5fd2fb284a7d470998af5083beb/ utils/compgen.py (ReCOGS) against ground truth.
We also measure grammar coverage (Zeller et al., 2023) (more detail in Appendix 9.10) by input examples supported by our RASP model against the full grammar of COGS/ReCOGS input sentences provided in the utilities of the IBM CPG project (Klinger et al., 2024) [^22] https://github.com/IBM/cpg/blob/ c3626b4e03bfc681be2c2a5b23da0b48abe6f570 /src/model/cogs_data.py#L523. See "Appendix: Methods Detail" (9.5). See also Appendix (9.1) for notebooks with results and steps to reproduce.
6 Results
Restricted Access Sequence Processing - grammar coverage using a flat pattern matching approach (not tree-based and not recursive) and autoregressive decoder loop
Figures 2, 5, 6, 7, 8, 9, 10, 11, 12, Table 1. See also Appendix (9.11) for more details. We selected 19 training-example equivalent sentences (identical to actual training examples after embedding to part-of-speech/verb-type, replacing with such training examples does not change the result) which cover 100% of the COGS non-recursive input grammar [^23] ReCOGS has same input grammar/vocabulary as COGS(lexical differences ignored, under the context-free grammar, tree-based assumption which is violated for our non-tree non-recursive model for prepositional phrases, see below for how that is handled) per (Zeller et al., 2023). Attention-head compatible flat patterns/rules were derived from those examples (Tables 1 , 7)242424see ”Appendix: Restricted Access Sequence Processing word-level token program/model design” (9.2). To handle prepositional phrases (recursive aspect of the input grammar) in a flat solution, we find it necessary on the training data [^25] E.g. center embedded pp example ”Isabella forwarded a box on a tree to Emma” where masking the center-embedded pp (”on a tree”) transforms the sentence to “Isabella forwarded a box _ to Emma” allows us to have noun-verb relationships extracted the same as other non-pp modified v_dat_p1 examples in Tables 1 , 7 by simply mapping the 3 nouns (Isabella, box, Emma) to agent, theme, recipient in order (same for all v_dat_p1). to add a rule that ignores "det common_noun" or "proper noun" preceded by a preposition when searching for noun indexes to report in relationships (agent, theme, recipient, etc) and as if we did that during pattern matching by using before/after matches instead of strict relative indexing (Figures 8 (ReCOGS) 11 (COGS), which also avoids "attraction errors", Figure 13). Sentential complements (recursive) in training also required masking (Figures 9, 12). The resulting 19 matchers implemented with attention-head compatible operations in RASP are given as part of speech / verb type sequences in Table 1 .
Restricted Access Sequence Processing - test set and generalization set performance on (Re)COGS
See ReCOGS results in Table 2 and COGS results in Table 4 .
(Wu et al., 2024) Encoder-Decoder Transformer from scratch baselines (ReCOGS_pos)
See Table 3.
(Wu et al., 2024) Encoder-Decoder baseline 2-layer Transformer does not improve on the obj_pp_to_subj_pp split when adding 1 or 2 additional layers (even allowing parameter count to increase) [^26] Since no improvement was observed, we did not run the costly experiments to increase the layers while controlling the parameter count (to distinguish improvements from increasing depth vs merely increasing the parameter count).
See Table 3.
Attraction Error Analysis for (Wu et al., 2024) baseline Encoder-Decoder Transformer on obj_pp_to_subj_pp split
(For additional methods detail see Appendix (9.8).) Of the obj_pp_to_subj_pp split single part errors in single verb sentences made by the (Wu et al., 2024) baseline EncDec Transformer where the agent was to the left of the verb [^27] Our hypothesis is in terms of nouns with a logical form relationship to a verb or other noun, where the relationship could be agent, theme, or recipient. Since the obj_pp_to_subj_pp split is in terms of subject vs object prepositional modification (instead of agent, recipient, or theme), we use the subset of sentences within this split where the agent is to the left of the verb and pp-modified as it corresponds to the subject in that case. , across n=10 models, 765 out of 767 (99.74%; 95% CI 99.06 to 99.97%) were in the agent part of the logical form (the predicted position for the error).
Critically across all n=10 (Wu et al., 2024) models, for 96.73% (740 of the 765 above; 95% CI (Clopper-Pearson) 95.21 to 97.87%) of the single point errors in logical forms for single verb sentences where the agent was on the left, modified by a prepositional phrase, and the error was in the agent part, the error in the logical form was that the agent index was accidentally assigned to the specific expected prepositional phrase noun (the one closest to the verb on the left side) instead of the original agent noun. (Figure 13) This does not vary much from randomly initialized and trained model to model, with the model-level average at 97.07% of such errors exactly as predicted (stderr=2.23% (n=10)), with 7 of 10 models having 100% of these errors exactly as predicted by our hypothesis [^28] Fraction for each model as predicted: [0.970, 0.761, 1.0, 1.0, 1.0, 1.0, 1.0, 1.0, 0.976, 1.0]). . The attraction to the nearest noun hypothesis predicts that the offset in the agent index varies with prepositional phrase recursion depth (as at depth > 1, there are multiple attractor prepositional nouns to choose from).  [^29] See ”Appendix: Attraction errors” (9.6) for examples of different pp recursion depths.
We report that for all (n=22) single logical form part errors observed (n=10 models, random weight init, random train data order, eval on n=1000 examples, combine errors) where in the input the agent is left of the verb and has a depth=2 prepositional phrase modification in this split, in 100% (95% CI (Clopper-Pearson) 84.6 to 100%; n=22) of those sentences the error in the agent right-index matched our prediction30303050% rate expected under null hypothesis of random pp noun substituting as agent rejected with p¡0.05 by Clopper-Pearson.
(Wu et al., 2024) Encoder-Decoder Transformer on new v_dat_p2 pp moved to recipient (from theme) split - as hard as hardest previous generalization split
See Figure 3 . Additional methods in Appendix 9.9. The baseline (Wu et al., 2024) EncDec Transformer was only able to achieve a Semantic Exact Match (sample mean +/- sample std) of 13% +/- 15.6% (n=10, random initial weights and train data order) with a 95% sample mean CI of 4% to 23%. Rejecting the null hypothesis of a parse tree-structured representation and instead consistent with the flat/non-hierarchical hypothesis (which does predict any center embedded pp in a novel position will cause errors) this new split v_dat_p2_pp_moved_to_recipient is as difficult or perhaps more difficult than the previous reported "hardest split" obj_pp_to_subj_pp .
(Wu et al., 2024) Encoder-Decoder Transformer trained with data augmented with v_dat_p2 pp moved to recipient (from theme) does NOT improve obj_pp_to_subj_pp performance
(Wu et al., 2024)’s baseline EncDec Transformer was trained with default data (ReCOGS_pos train.tsv) but with additionally the same modified v_dat_p2 pp training examples used for the "v_dat_p2_pp_moved_to_recipient" split (non-subject recipient modified with prepositional phrase, so nonoverlapping with subj_pp, Figure 3 (b) and (d)) above on which it performed poorly, then tested on the standard prepositional modification generalization split "obj_pp_to_subj_pp", after which it achieved 22% +/- 6.7% Semantic Exact Match (sample mean +/- std, n=10) with 95% sample mean CI of 17.9% to 26.1% (not significantly different than (Wu et al., 2024)’s baseline by one-tailed Welch’s unequal variances t-test).
7 Analysis
Our RASP model of a Transformer Encoder-Decoder (EncDec), without using tree-structured representations (nor recursive tree-structured rules like ‘np_det pp np -> np_pp -> np‘) attains near perfect scores on the (Re)COGS generalization splits, even the structural generalizations (Tables 2 and 4). This includes prepositional phrase recursion and sentential complement recursion generalization splits up to depth 12 based on rules that can be derived from the depth 2 training examples, without any hardcoded prepositional phrase or sentential complement expansion shortcuts added313131a single rule applies to all depths; the only limit on length generalization is the RASP interpreter and a simple to extend positional encoding which only handles sentences up to a limit.
Thus, we demonstrated by construction using the Restricted Access Sequence Processing language which can be compiled to concrete Transformer weights that theoretically a Transformer EncDec can solve the COGS input to COGS/ReCOGS_pos logical form translation in a systematic, compositional, and length generalizing way.
Recall we found a single flat masking rule we originally added to fit center embedded pp training examples, to ignore "pp det common_noun" and "pp proper_noun" when matching nouns for the agent, theme, recipient right indices, was sufficient to avoid structural generalization errors due to pp modification in novel positions.
Interestingly, we imagined ablating that single rule and hypothesized attraction to the nearest noun (Figure 13 and "Appendix: Attraction Errors" (9.6)) in its absence and found this predicted the exact error (the nearest noun to the verb on the expected side is mistaken for the agent of the verb) in 96% of the single relationship errors the (Wu et al., 2024) baseline Transformers make on the obj-pp-to-subj-pp split323232overall, their semantic exact match on the split is measured by us at 19.7%, consistent with their Figure 5 when the agent is left of the verb in single verb sentences (suggesting perhaps that the baseline (Wu et al., 2024) Transformer trained from scratch is also not learning a hierarchical, tree-structured representation.)333333(Li et al., 2023) also observe this stating ”For instance, in sentences like ’A cat on the mat froze’, models often misinterpret the closer NP the mat as the subject.” Our explanation could have been refuted by other single relationship errors occurring as frequently as the agent, indicating general model confusion (independently getting incorrect agent and theme, not just agent relationships) and/or when making an agent error, the model could have simply put nonsense indices or referred to any other position other than the closest pp noun position to the verb (which does vary and depends on pp depth) to refute our hypothesis.
The flat/non-tree structured hypothesis for the baseline Transformer can also be checked by making a prediction on a completely different syntax, given we predict any novel center pp substitution will be challenging (not about subjects vs objects): the ‘np v_dat_p2 np pp np np‘ prepositional phrase modification (where the recipient relationship is modified instead of the subject/agent so is a distinct check of our hypothesis) [^34] See Appendix 9.9 and we indeed found that this was at least as hard as the most difficult split previously reported, the ‘obj_pp_to_subj_pp‘ split (Figure 3).
Figure 3: (Wu et al., 2024) Encoder-Decoder Transformer trained from scratch generalizing to new v_dat_p2 pp moved to recipient (from theme) split is as hard as the previously reported hardest generalization split consistent with the flat/non-recursive/non-tree representation hypothesis ((d) rejects $H_{0}$). Figure 5 shows how a flat RASP model can recognize (b).
Maybe (Wu et al., 2024) baseline EncDec is depth-constrained to learn a non-hierarchical, non-tree model with these characteristic errors but with more layers it would learn to recursively combine ‘np_det pp np -> np_pp -> np‘ (to some fixed depth, limited by the number of Transformer blocks) and perform better on pp related splits353535(Csordás et al., 2022): ”the network should be sufficiently deep, at least as deep as the deepest data dependency in the computational graph built from elementary operations (e.g., in the case of a parse tree, this is the depth of the tree)”. But the hierarchical approach may not be a scalable approach, as we must make the network ever deeper to handle increased pp depths instead of just looping (without weight-sharing), and deeper parses encountered at inference time than seen in training would activate untrained layers, leading to generalization failures. As shown by the RASP model, flat, but with pp masking (Figure 8 (ReCOGS) 11 (COGS)), may be a better strategy.. However, training a (Wu et al., 2024) baseline EncDec Transformer from scratch we found no benefit to 3-4 layers instead of 2 on the ReCOGS obj_pp_to_subj_pp split, consistent with (Petty et al., 2024)’s finding on COGS.
These results suggest we may interpret the poor performance on generalizing on unseen pp related modification related splits as arising from the baseline 2-4 layer EncDec Transformers learning a non-tree representation that cannot leverage the recursive grammar rule ‘np_det pp np -> np_pp -> np‘ during learning, thus requiring them to actually observe more of the various pp substitutions to learn them [^36] See ”Appendix: Composition and Learning” (9.13) (if pp masking as used by our RASP model is not learned instead).
8 Conclusion
Using Restricted Access Sequence Processing immediately helped us discover additional failure modes (e.g. new "v_dat_p2_pp_moved_to_recipient" split [^37] After this paper was written we found our predicted split ”v_dat_p2_pp_moved_to_recipient” has also been added to an extended separate (SLOG) version of COGS in (Li et al., 2023), sec 2.2.1 indirect object mod (4), confirmed difficult ) of (Wu et al., 2024)’s baseline EncDec Transformer, predict logical form errors in detail, and demonstrated that tree-structured representations are not needed to cover the recursive aspects of the (Re)COGS grammar. We recommend RASP for understanding Transformer behavior even for tasks as complex as COGS. We predict that Transformers should be able to perform the (Re)COGS tasks accurately for novel prepositional phrase substitution positions or structural recursion depths unseen in training and that the problem is just of getting the Transformer to learn the appropriate rules383838e.g. to ignore ”pp det common_noun” and ”pp proper_noun” when finding nouns in relationships with verbs, which allows our RASP models to attain near perfect scores on (Re)COGS structural and depth generalizations, turning attention to data augmentation [^39] We tried one augmentation (A9.9) out of many possible. (Qiu et al., 2022)404040contrasting (Qiu et al., 2022), the RASP-for-(Re)COGS flat solutions generalize to arbitrary depth (not just the depth of recursion demonstrated in augmented data), curriculum learning (Bengio et al., 2009), reinforcement learning (Ranzato et al., 2016), training objectives (Ahuja et al., 2024), and other approaches (Csordás et al., 2021) (Ontanon et al., 2022).
Known Project Limitations
The Restricted Access Sequence Processing code is not optimized [^41] A slow RASP program does NOT necessarily imply the equivalent Transformer would be slow.. Cannot yet predict attention heads and layers required from the select and aggregate operations performed like the RASP authors (Weiss et al., 2021) were able to do with their problems.
Grammar coverage (Zeller et al., 2023) is only valid when the expansions are rules your model can learn. [^42] If for example, as with our flat RASP model by design or as we hypothesize for (Wu et al., 2024)’s baseline Encoder-Decoder Transformer, the model cannot or will not learn the rule ‘np_det pp np -¿ np_pp -¿ np‘ which recursively replaces noun phrases modified by prepositional phrases with a noun phrase, then grammar coverage will assume any prepositional phrase exposure is sufficient, which is evidently not true given the errors on prepositional phrase modification generalization splits reported here and by (Wu et al., 2024), (Kim and Linzen, 2020). We specifically made use of this limitation in this paper but still caution anyone about it who might just take the grammar coverage metric away by itself.
The attraction error analysis of the (Wu et al., 2024) baseline Encoder-Decoder Transformer on the obj_pp_to_subj_pp split does not yet attempt to explain the common case of multiple errors in the logical form.434343e.g. agent index may be replaced by prepositional phrase noun but also a spurious theme relationship is added or the theme index is also corrupted
Much deeper Transformer networks may be learning a tree-based grammar representation [^44] Nothing explored here rules that out and there is plenty of evidence outside the COGS task-related literature suggesting this will be the case: (Tenney et al., 2019) show the 24-layer BERT model seems to handle ”POS tagging, parsing, NER, semantic roles, then coreference”; (Hewitt and Manning, 2019) ”provid[e] evidence that entire syntax trees are embedded implicitly in deep models’ [including BERT’s] vector geometry”, and (Goldberg, 2019) shows BERT excels at subject-verb agreement, ”which [is] traditionally taken as evidence for the existence [of] hierarchical structure” (though e.g. in this work we see that ignoring distractor nouns in long-term dependencies does not require hierarchy or a deep understanding of syntax but simple rules like ignore ”pp det common_noun” and ”pp proper_noun” for finding noun-verb relationships in the logical form can allow for handling of such long-range dependencies). On the other hand, (Petty et al., 2024) argue specifically for the COGS benchmark (semantically equivalent to ReCOGS which is derived from it) that increasing depth does not allow their Transformers to make progress on the structural generalization splits, even at depths up to 32 layers. and not suffer from the predicted generalization issues observed in (Wu et al., 2024)’s baseline 2-layer Transformer and predicted by our intentionally non-tree RASP model (if compensating rules to avoid attraction errors in a flat model are not also learned).
As stated elsewhere in the paper, we are focused on the structural generalizations that Transformers currently struggle with (prepositional phrase recursion (pp recursion), sentential complement (cp) recursion, object-pp to subj-pp modification generalization (obj-pp-to-subj-pp)) and not lexical generalizations (which Transformers are demonstrated to do well on by others). Thus, as noted elsewhere, our model assumes the embedding can map words to their possible part-of-speech and verb-type and does not address the learnability of this mapping (number of exposures required, etc).
The (Re)COGS tasks are English specific and our findings on compositional generalization may not necessarily apply (or apply differently) to other languages. Our description of "attraction errors" for example does not appear specific to subject-verb-object ordering in English but a similar analysis should be conducted in other languages and we used the obj-pp-to-subj-pp split in our detailed error analysis which is subject-verb-object ordering sensitive.
The RASP-for-(Re)COGS models applied to unintended use WILL give invalid results or halt - we have NOT provided a general language model, we have provided a simulation of how a Transformer could perform a specific task. The RASP-for-(Re)COGS models/simulations as provided are for research purposes only to prove feasibility of the (Re)COGS tasks by Transformers and is not appropriate for ANY other uses whatsoever without modification. For one, an actual Transformer performing the equivalent operations would run orders of magnitude faster, which should be reason enough to not want to use the RASP simulation for actual input-output tasks outside of a research setting. However, there is also no "tokenizer" provided for the RASP model to handle out-of-vocabulary inputs and fallback paths for out-of-grammar examples are not provided so the RASP model will halt on most inputs and can only run on the in-distribution (non-augmented) training data, and the dev, test, and gen sets of (Re)COGS, though such aspects could be added. We provide the code for reproducing the results of this study and for researchers who are capable of writing RASP themselves to build upon the work and/or more easily apply RASP to their own problems given our examples, not for immediate application to any other tasks without appropriate modification.
See also "Appendix: Potential Risks" (9.14) and "Appendix: Scientific Artifacts - Is Our Use Consistent with Authors’ Intention and Licensing" (9.15) for other potential concerns.